OpenStudy (anonymous):

GIVE MEDAL AND BECOME FAN!!! In ∆ABC, BC = 4 centimeters, m B = m C, and m A = 20°. What is AC to two decimal places? 9.32 centimeters 10.01 centimeters 11.52 centimeters 12.09 centimeters

OpenStudy (anonymous):

OMG love ur hair!!!

OpenStudy (misty1212):



OpenStudy (misty1212):

HI!! thanks !

OpenStudy (misty1212):

does that picture look right?

OpenStudy (anonymous):

yes but its confusing

OpenStudy (misty1212):

those two bottom angles are equal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im fanning you :)

OpenStudy (misty1212):

no that is wrong, each are 80

OpenStudy (misty1212):

since \[80+80+20=180\]

OpenStudy (misty1212):

you good with that? now it is time for the law of sines

OpenStudy (anonymous):

yea i did and im ready >:)

OpenStudy (misty1212):

\[\frac{AC}{\sin(80)}=\frac{4}{\sin(20)}\]

OpenStudy (misty1212):

so \[AC=\frac{4\sin(80)}{\sin(20)}\]

OpenStudy (anonymous):

do i divide tht?

OpenStudy (misty1212):

use this http://www.wolframalpha.com/input/?i=4sin%2880%29%2Fsin%2820%29

OpenStudy (misty1212):

you need a calculator to do it anyway might as well use wolfram if you scroll down you will see the answer is about \(11.52\)

OpenStudy (anonymous):

thx girl :)

OpenStudy (misty1212):

\[\huge \color\magenta\heartsuit\]